u+v =3,4 and u ➖ v=1, ➖ 2 then find v and u

1
Expert's answer
2022-06-07T08:19:52-0400
"u_1+v_1=3"

"u_1-v_1=1"

Then "u_1=2, v_1=1."


"u_2+v_2=4"

"u_2-v_2=2"

Then "u_2=3, v_2=1."


Therefore "\vec{u}=\langle2, 3\rangle,\vec{v}=\langle1, 1\rangle."